# short-time Fourier transform iSTFT inverse short-time Fourier transform

### v1

@@ III-A Diffusion-Based Generative Models
parameters $\psi$, speech estimate $\hat{\mathbf{x}}_{0}$
@@
$\mathbf{A}\leftarrow\exp{\mathrm{lerp}(\log{\mathbf{A}^{\prime}})}$ ) $\triangleright$ Frequency interpolation
$\mathbf{H}=\mathbf{A}e^{j\mathbf{\Phi}}$
$\overline{\mathbf{H}}=\mathrm{STFT}(\delta_{\text{d}}\oplus\mathcal{P}_{\text{min}}(\mathrm{iSTFT}(\mathbf{H})))$triangleright Projection step
$\hat{\mathbf{X}}=\mathrm{STFT}(\hat{\mathbf{x}}_{0})$
@@
III-B Diffusion Posterior Sampling for Dereverberation
We discuss in this section how diffusion priors can be adapted in order to solve inverse problems. While some traditional methods derive maximum a posteriori estimators for blind dereverberation [9, 10, 11], we exploit the generative nature of diffusion models to solve this inverse problem using posterior sampling. Assuming for now that the RIR $\mathbf{h}$ is known, we attempt to sample from the posterior distribution of the anechoic speech given the measurement and the RIR $p(\mathbf{x}_{0}|\mathbf{y},\mathbf{h})$. This is achieved by solving the probability flow ODE (1), replacing the score function $\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{x}_{\tau})$ by the posterior score $\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{x}_{\tau}|\mathbf{y},\mathbf{h})$ [23]. The posterior score is obtained through Bayes’ rule as:
$\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{x}_{\tau}|\mathbf{y},\mathbf{h})=\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{x}_{\tau})+\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{y}|\mathbf{x}_{\tau},\mathbf{h})\,.$
The first term, or prior score, is directly obtained via the score model $\mathbf{s}_{\theta}(\mathbf{x}_{\tau},\tau)$. The likelihood score $\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{y}|\mathbf{x}_{\tau},\mathbf{h})$ is in general intractable for $\tau>0$. As in [27], we employ an estimate of $\mathbf{x}_{0}$, denoted as $\hat{\mathbf{x}}_{0}$, and we assume that this estimate is a sufficient statistic for $\mathbf{x}_{\tau}$. This results in a first assumption $p(\mathbf{y}|\mathbf{x}_{\tau},\mathbf{h})\approx p(\mathbf{y}|\hat{\mathbf{x}}_{0},\mathbf{h})$. The estimate $\hat{\mathbf{x}}_{0}(\mathbf{x}_{\tau})$ is obtained as the posterior mean of $\mathbf{x}_{0}$ knowing $\mathbf{x}_{\tau}$ and is derived using Tweedie’s formula, i.e. one-step denoising of $\mathbf{x}_{\tau}$:
$\hat{\mathbf{x}}_{0}(\mathbf{x}_{\tau})\overset{\Delta}{=}\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{\tau}]\approx\mathbf{x}_{\tau}-\sigma^{2}(\tau)\mathbf{s}_{\theta}(\mathbf{x}_{\tau},\tau).$ (6)
In order to approximate $p(\mathbf{y}|\hat{\mathbf{x}}_{0},\mathbf{h})$, previous work [26] models the error between $\mathbf{y}$ and $\mathbf{h}\ast\hat{\mathbf{x}}_{0}$ to follow a zero-mean Gaussian distribution in the time domain. The corresponding expression for the likelihood score $\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{y}|\mathbf{x}_{\tau},\mathbf{h})$ is then a simple weighted $L^{2}$-distance between $\mathbf{y}$ and $\mathbf{h}\ast\hat{\mathbf{x}}_{0}$. However we observed that far better dereverberation performance and speech quality can be achieved by substituting the obtained distance with a $L^{2}$-distance between compressed spectrograms instead. This is analogous to modelling the likelihood score as:
$\nabla_{\mathbf{x}_{\tau}}\log p(\mathbf{y}|\mathbf{x}_{\tau},\mathbf{h})\approx-\zeta(\tau)\nabla_{\mathbf{x}_{\tau}}\mathcal{C}(\mathbf{y},\mathbf{h}\ast\hat{\mathbf{x}}_{0}(\mathbf{x}_{\tau}))\,,$ (7)
where $\zeta(\tau)$ is a diffusion-time-dependent scaling parameter that controls the influence of the likelihood score term in the sampling trajectory, and $\mathcal{C}(\cdot,\cdot)$ is the following cost function:
$\mathcal{C}(\mathbf{u},\mathbf{v})=\frac{1}{M}\sum_{m=1}^{M}\sum_{k=1}^{K}\rVert S_{\text{comp}}(\mathbf{u})_{m,k}-S_{\text{comp}}(\mathbf{v})_{m,k}\lVert_{2}^{2}.$ (8)
There, $S_{\text{comp}}(\mathbf{u})=|\text{STFT}(\mathbf{u})|^{2/3}\exp{j\angle\text{STFT}(\mathbf{u})}$ ) is the magnitude-compressed spectrogram of $\mathbf{u}$. This cost function implies that we model the reconstruction error in the compressed STFT domain as a Gaussian with unknown variance $\frac{1}{2\zeta(\tau)}$. We apply this compression to boost low-energy components as typically observed in high frequencies of speech signals or in late reverberation tails, and account for the heavy-tailedness of speech distributions [4]. Such a strategy is also employed in [46] for data representation.
$\zeta(\tau)=\frac{\sqrt{L}\,\tilde{\zeta}}{\norm{\nabla_{\mathbf{x}_{\tau}}\mathcal{C}\left(\mathbf{y},\mathbf{h}\ast\hat{\mathbf{x}}_{0}(\mathbf{x}_{\tau})\right)}_{2}}\,,$ (9)
where $\tilde{\zeta}$ is a fixed coefficient.
IV Blind Diffusion-Based Dereverberation
@@
     $\hat{\mathbf{x}}_{0}\leftarrow\mathrm{Rescale}(\hat{\mathbf{x}}_{0})$triangleright Constrain RMS power
@@
     $\mathbf{g}_{n}\leftarrow\zeta(\tau_{n})\nabla_{\mathbf{x}_{n}}\mathcal{C}(\mathbf{y},\mathcal{A}_{\psi_{n-1}}(\hat{\mathbf{x}}_{0}))$triangleright LH score approx.
@@
$\mathbf{x_{\text{init}}}$$\sigma_{N}\mathbf{z}$$\mathbf{y}$
@@
This section elaborates on the proposed method for blind dereverberation, where the impulse response $\mathbf{h}$ is unknown. In Section IV-A, we define a reverberation operator $\mathcal{A}_{\psi}(\cdot)$, which comprises a structured parametric model of the RIR, with parameters $\psi$. This operator is summarized in Algorithm 1. The proposed inference method BUDDy, detailed in IV-B, performs joint speech dereverberation and RIR estimation by combining the conditional sampling from a diffusion model with an optimization of the RIR parameters. The complete inference procedure is summarized in Algorithm 2, and the processing pipeline is visualized in Fig.2.
IV-A Reverberation Operator
IV-A1 Subband Filtering
In contrast to Section III, here we model reverberation using a subband filtering approximation in the short-time Fourier transform (STFT) domain [48, 49]. Let $\mathbf{H}:=\mathrm{STFT}(\mathbf{h})\in\mathbb{C}^{N_{\mathbf{h}}\times K}$ represent the STFT of a RIR $\mathbf{h}$ with $N_{\mathbf{h}}$ time frames and $K$ frequency bins. Similarly, let $\mathbf{X}\in\mathbb{C}^{M\times K}$, and $\mathbf{Y}$, denote the STFTs of anechoic $\mathbf{x}_{0}$ and reverberant $\mathbf{y}$ speech signals, repectively. The subband convolution operation applies independent convolutions along the time dimension of each frequency band:
@@ IV-A1 Subband Filtering
The resulting reverberant signal $\mathbf{Y}\in\mathbb{C}^{(M+N_{\mathbf{h}}-1)\times K}$ can be transformed to time domain by applying the inverse STFT. The subband filtering model only approximates the time-domain convolution, as it does not account for the spectral leakage between frequency bands. However, it is empirically found to be a valid assumption in many scenarios involving reverberation [49, 7, 50]. In our case, we noticed that adding 50% zero-padding to the end of the frames before computing the STFT was important to avoid cyclic convolution artifacts when retrieving the resulting signal to the time domain.
IV-A2 Room Impulse Response Prior
In the blind scenario, we need to estimate $\mathbf{H}$, which is an ill-posed problem task when not knowing the anechoic speech. Therefore, we need to constrain the space of possible solutions by imposing a prior on $\mathbf{H}$. We propose a structured, differentiable prior on $\mathbf{H}$, whose parameters $\psi$ can be estimated through gradient descent. We denote the complete forward reverberation operator, including forward and inverse STFT operations, as $\mathcal{A}_{\psi}(\cdot):\mathbb{R}^{L}\rightarrow\mathbb{R}^{L+L_{\mathbf{h}}-1}$. The whole processing pipeline is summarized in Algorithm 1 with each component detailed below.
@@ IV-A2 Room Impulse Response Prior
Note that our parameterization can be extended to model coupled spaces by employing several decay parameters per band and summing their respective contributions [51]. We found it beneficial to constrain $w_{b}$ and $\alpha_{b}$ within a limited range to stabilize the optimization, specially at early stages. This is achieved by clamping the parameters to predefined minimum and maximum values after every optimization iteration, as specified in Appendix -C1. Once the parameters are estimated, we reconstruct the $K$-bands magnitudes $\mathbf{A}$ by interpolating the subsampled $\mathbf{A}^{\prime}$ as $\mathbf{A}=\exp(\mathrm{lerp}(\log(\mathbf{A}^{\prime})))$, where $\mathrm{lerp}$ represents linear interpolation on the frequency scale. For this purpose, we employ the $\mathrm{torchcde}$ library, which facilitates efficient and differentiable interpolation [52]. After interpolation of the magnitude matrix, we then obtain the time-frequency RIR $\mathbf{H}$ by multiplying the magnitude matrix $\mathbf{A}$ with the complex phase exponentials:
@@
IV-A3 Projections
We extend our forward reverberation operator with a series of projections to increase the likelihood of generating plausible RIRs. Thus, the time-frequency RIR $\mathbf{H}$ is further processed as:
@@ IV-A3 Projections
This primarily ensures STFT consistency of $\overline{\mathbf{H}}$, exploiting the redundancy of the STFT representation and imposing inter-frame correlations between the RIR phases $\Phi$. We then enforce that the time-domain RIR estimate $\mathbf{h}$ has minimum-phase lag, using the Hilbert transform-based method in [53]. This is indicated by the operator $\mathcal{P}_{\text{min}}$ and guarantees stability of the inverse RIR filter [2]. We refer the reader to Appendix -A for further details. Finally, the operation $\delta\oplus(\cdot)$ replaces the first sample of the time-domain RIR with a unit impulse. This has the effect of injecting knowledge of the direct path in $\overline{\mathbf{H}}$, and further requires us to correct the magnitude matrix $\mathbf{A}$ to account for this operation. It is important to note that these steps are integral to the reverberation operator $\mathcal{A}_{\psi}(\cdot)$, which maps the parameters $\psi$ to the convolved signal $\mathcal{A}_{\psi}(\hat{\mathbf{x}}_{0})$, as outlined in Algorithm 1. Since all operations are differentiable, we compute gradients with respect to $\psi$ by backpropagating through all operations. We propose a detailed ablation study of these projection and correction steps in Section V-A6.
IV-B Blind Dereverberation Inference
We aim to solve the following joint dereverberation and RIR parameter optimization problem:
$\hat{\mathbf{x}}_{0},\hat{\psi}=\underset{\mathbf{x}_{0},\psi}{\mathrm{arg\,min}}\;\mathcal{C}(\mathbf{y},\mathcal{A}_{\psi}(\mathbf{x}_{0}))+\mathcal{R}(\psi)\;\;\text{s.t.}\;\;\mathbf{x}_{0}\sim p_{\text{data}}$ (14)
where $\mathcal{C}(\mathbf{y},\mathcal{A}_{\psi}(\mathbf{x}_{0}))$ is the reconstruction error with $\mathcal{C}$ the cost function introduced in (8), and $\mathcal{R}(\psi)$ is a RIR prior regularization term. This objective seeks to find the optimal speech $\hat{\mathbf{x}}_{0}$ and RIR parameters $\hat{\psi}$ that minimize both losses while imposing the soft constraint that the estimated signal $\hat{\mathbf{x}}_{0}$ should adhere to the anechoic speech distribution $p_{\text{data}}$. We leverage the pre-trained score model $\mathbf{s}_{\theta}(\mathbf{x}_{\tau},\tau)$ trained on anechoic speech to enforce this constraint.
We optimize (14) by solving the following ODE, obtained from the classical probability-flow ODE (1) where we injected the specified diffusion parameters of Karras et al. [44] and the likelihood score approximation derived in (7):
$\mathrm{d}\mathbf{x}_{\tau}=-\tau\left[\mathbf{s}_{\theta}(\mathbf{x}_{\tau},\tau)-\zeta(\tau)\displaystyle\nabla_{\mathbf{x}_{\tau}}\mathcal{C}(\mathbf{y},\mathcal{A}_{\psi_{\tau}}(\hat{\mathbf{x}}_{0}))\right]\mathrm{d}\tau\,.$ (15)
The regularization term $\mathcal{R}(\psi)$ introduced in (14) is:
$\mathcal{R}(\psi)=\frac{1}{N_{\mathbf{h}}}\sum_{m=1}^{N_{\mathbf{h}}}\sum_{k=1}^{K}\rVert S_{\text{comp}}(\hat{\mathbf{h}}_{\psi})_{m,k}-S_{\text{comp}}(\hat{\mathbf{h}}_{\psi^{\prime}}+\sigma^{\prime}\mathbf{v})_{m,k}\lVert_{2}^{2},$ (16)
where $\hat{\mathbf{h}}_{\psi}=\mathcal{A}_{\psi}(\delta)$ is the current time-domain RIR estimate and $\mathbf{v}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is a vector of white Gaussian noise. The noise level schedule $\sigma^{\prime}(\tau)$ is similar to the score model schedule $\sigma(\tau)$ but its values are limited to avoid having too much noise at early steps (see Appendix -C1) The term $\hat{\mathbf{h}}_{\psi^{\prime}}$ represents a copy of $\hat{\mathbf{h}}_{\psi}$, such that the $\mathrm{arg\,min}$ in (14) does not apply to it.. In other terms, we detach the gradients of $\hat{\mathbf{h}}_{\psi}$ from the optimization graph to obtain $\hat{\mathbf{h}}_{\psi^{\prime}}$, similar to what is done in e.g. [54] for guiding data reconstruction. We provide in Appendix -B a short analysis of the regularization objective $\mathcal{R}(\psi)$. We show that this term injects multiplicative noise with standard deviation $\sigma^{\prime}(\tau)$ in the RIR parameter gradients, which arguably smoothes the RIR parameter search.
During optimization, we further rescale the denoised speech estimate $\hat{\mathbf{x}}_{0}$ so that its root-mean-square power (RMS) matches the average RMS power of clean speech computed on the training set. Using this additional constraint helps lifting the indeterminacy when jointly optimizing the speech $\mathbf{x}_{0}$ and RIR parameters $\psi$. This stesp is included in our ablation study in Section V-A6.
@@ IV-B Blind Dereverberation Inference
In order to guide and accelerate reverse diffusion, it is often beneficial to use warm initialization, i.e. to let the reverse diffusion process start from a speech sample $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{x}_{\text{init}},\sigma^{2}(T)\mathbf{I})$, where $\mathbf{x}_{\text{init}}$ has some interesting cues about the clean signal we wish to estimate. Similar to [34], we obtain the initial mean signal $\mathbf{x}_{\text{init}}$ through WPE [7], a blind dereverberation algorithm based on variance-normalized delayed linear prediction. WPE performs mild dereverberation, which allows us to get closer to the clean speech, while not introducing too many distortions to the signal. As WPE is blind and unsupervised, our method remains fully blind and unsupervised as well.
V Experiments and Results

### v2

@@ III Blind Diffusion-Based Dereverberation
function $\mathcal{A}_{\psi}$($\hat{\mathbf{x}}_{0}$)
@@
     $\mathbf{A}\leftarrow\exp{\left(\mathrm{lerp}(\log{\mathbf{A}^{\prime}})\right)}$ ) $\triangleright$ Frequency interpolation
     $\mathbf{H}\leftarrow\mathbf{A}\cdot e^{j\mathbf{\Phi}}$
     $\overline{\mathbf{H}}\leftarrow\mathrm{STFT}(\delta_{\text{d}}\oplus\mathcal{P}_{\text{min}}(\mathrm{iSTFT}(\mathbf{H})))$triangleright Projection step
     $\hat{\mathbf{X}}\leftarrow\mathrm{STFT}(\hat{\mathbf{x}}_{0})$
@@
     $\hat{\mathbf{x}}_{0}\leftarrow\mathrm{Rescale}(\hat{\mathbf{x}}_{0})$triangleright Constraint RMS power
     $\mathbf{g}_{n}\leftarrow-\zeta(\tau_{n})\nabla_{\mathbf{x}_{n}}\mathcal{C}(\mathbf{y},\mathcal{A}_{\psi_{n}}(\hat{\mathbf{x}}_{0}))$triangleright LH score approx.
   Phase 2 – M-step
@@
This section elaborates on the proposed method BUDDy, which extends the informed method presented in prior work [26] to the blind scenario, where the impulse response $\mathbf{h}$ is unknown. In Section III-A, we define a reverberation operator $\mathcal{A}_{\psi}(\cdot)$, which comprises a structured parametric model of the RIR, with parameters $\psi$. Section III-B then details the posterior sampling scheme used to obtain a speech utterance from the estimated posterior distribution. Finally, Section III-C describes the optimization procedure for the reverberation model parameters $\psi$ using an expectation-maximization (EM) formulation. The complete inference procedure is summarized in Algorithm 2, and an overview of the processing pipeline is given in Fig. 1.
III-A Reverberation Operator
III-A1 Subband Filtering
In contrast to [26] where the original time-domain convolution model is adopted, here we model reverberation using a subband filtering approximation in the STFT domain [50, 51]. This approach enables us to incorporate prior knowledge about the characteristics of reverberation through a structured model of RIR magnitudes, characterized by exponential decays in each subband. By modeling reverberation in this way, we facilitate optimization and reduce the complexity of the RIR parameter search space. Let $\mathbf{H}:=\mathrm{STFT}(\mathbf{h})\in\mathbb{C}^{N_{\mathbf{h}}\times K}$ represent the STFT of a RIR $\mathbf{h}$ with $N_{\mathbf{h}}$ time frames and $K$ frequency bins. Similarly, let $\mathbf{X}\in\mathbb{C}^{M\times K}$, and $\mathbf{Y}$, denote the STFTs of anechoic $\mathbf{x}_{0}$ and reverberant $\mathbf{y}$ speech signals, repectively. The subband convolution operation applies independent convolutions along the time dimension of each frequency band:
@@ III-A1 Subband Filtering
The resulting reverberant signal $\mathbf{Y}\in\mathbb{C}^{(M+N_{\mathbf{h}}-1)\times K}$ can be transformed to time domain by applying the inverse STFT. The subband filtering model only approximates the time-domain convolution, as it does not account for the spectral leakage between frequency bands. However, it is empirically found to be a valid assumption in many scenarios involving reverberation [51, 7, 52]. Adding 50% zero-padding to the end of the frames before computing the STFT is important to avoid cyclic convolution artifacts when transforming the resulting signal back to the time domain.
III-A2 RIR Prior
In the blind scenario, estimating $\mathbf{H}$ is an ill-posed problem when the anechoic speech is unknown. Therefore, we need to constrain the space of possible solutions by imposing a prior on $\mathbf{H}$. We propose a structured, differentiable prior on $\mathbf{H}$, whose parameters $\psi$ can be estimated with gradient-based optimizers like Adam [53]. We denote the complete forward reverberation operator, including forward and inverse STFT operations, as $\mathcal{A}_{\psi}(\cdot):\mathbb{R}^{L}\rightarrow\mathbb{R}^{L+L_{\mathbf{h}}-1}$. The whole processing pipeline is summarized in Algorithm 1 with each component detailed below.
@@ III-A2 RIR Prior
Note that our parameterization can be extended to model coupled spaces by employing several decay parameters per band and summing their respective contributions [54]. Once the parameters are estimated, we reconstruct the $K$-bands magnitudes $\mathbf{A}$ by interpolating the subsampled matrix $\mathbf{A}^{\prime}$ as $\mathbf{A}=\exp(\mathrm{lerp}(\log(\mathbf{A}^{\prime})))$, where $\mathrm{lerp}$ represents linear interpolation on the frequency scale. For this purpose, we employ the $\mathrm{torchcde}$ library, which facilitates efficient and differentiable interpolation [55]. After interpolation of the magnitude matrix, we then obtain the time-frequency RIR $\mathbf{H}$ by multiplying the magnitude matrix $\mathbf{A}$ with the complex phase exponentials:
@@
III-A3 Projections
We extend our forward reverberation operator with a series of projections to increase the likelihood of generating plausible RIRs. Thus, the time-frequency RIR $\mathbf{H}$ is further processed as
@@ III-A3 Projections
This primarily ensures STFT consistency of $\overline{\mathbf{H}}$, exploiting the redundancy of the STFT representation and imposing inter-frame correlations between the RIR phases $\boldsymbol{\Phi}$. We then enforce that the time-domain RIR estimate $\mathbf{h}$ has minimum-phase lag, using the Hilbert transform-based method in [56]. This is indicated by the operator $\mathcal{P}_{\text{min}}$ and guarantees stability of the inverse RIR filter [2]. We refer the reader to Appendix -A for further details. Finally, the operation $\delta\oplus(\cdot)$ replaces the first sample of the time-domain RIR with a unit impulse. This has the effect of injecting knowledge of the direct path in $\overline{\mathbf{H}}$, and further requires us to correct the magnitude matrix $\mathbf{A}$ to account for this operation. It is important to note that these steps are integral to the reverberation operator $\mathcal{A}_{\psi}(\cdot)$, which maps the parameters $\psi$ to the convolved signal $\mathcal{A}_{\psi}(\hat{\mathbf{x}}_{0})$, as outlined in Algorithm 1. Since all operations are differentiable, we compute gradients with respect to $\psi$ by backpropagating through all operations. We propose a detailed ablation study of these projection and correction steps in Section IV-A6.
III-B Posterior Speech Sampling
For sampling a speech utterance from the posterior distribution $p_{\psi}(\mathbf{x}_{0}|\mathbf{y})$, we adapt the posterior sampling algorithm of Section II-B to the blind setting. As in Section II-B, the pre-trained score model $\mathbf{s}_{\theta^{\ast}}(\mathbf{x}_{\tau},\tau)$ is used to provide an implicit prior $p_{\theta^{\ast}}(\mathbf{x}_{0})$ on anechoic speech with fixed parameters $\theta^{\ast}$. The likelihood $p_{\psi}(\mathbf{y}|\mathbf{x}_{0})$ is approximated as
$p_{\psi}(\mathbf{y}|\mathbf{x}_{0})\propto\mathrm{exp}\left(-\zeta(\tau)\mathcal{C}(\mathbf{y},\mathcal{A}_{\psi}(\hat{\mathbf{x}}_{0}^{\tau})\right)\,,$ (16)
which is analogous to (8), but with the proposed reverberation operator $\mathcal{A}_{\psi}(\cdot)$ applied in place of a convolution with the oracle RIR, since the latter is unavailable in our blind scenario. The sampling procedure then follows the ODE (II-B) where the likelihood term $p_{\psi}(\mathbf{y}|\mathbf{x}_{0})$ defined in (16) is used instead of the RIR-informed likelihood $p(\mathbf{y}|\mathbf{x}_{0})$ from (8).
@@ III-B Posterior Speech Sampling
In order to guide and accelerate reverse diffusion, it is beneficial to use warm initialization, i.e., let the reverse diffusion process start from a speech sample $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{x}_{\text{init}},\sigma^{2}(T)\mathbf{I})$, where $\mathbf{x}_{\text{init}}$ contains relevant information about the clean signal we wish to estimate. Similar to [34], we obtain $\mathbf{x}_{\text{init}}$ through WPE [7], a blind dereverberation algorithm based on variance-normalized delayed linear prediction. WPE performs mild dereverberation, which allows us to get closer to the clean speech, while not introducing much distortion to the signal. As WPE is blind and unsupervised, our method remains fully blind and unsupervised as well.
III-C Reverberation Model Parameter Optimization
The posterior speech sampling approach described in Section III-B relies on approximating the likelihood $p_{\psi}(\mathbf{y}|\mathbf{x}_{0})$, where the main challenge is the dependence on the unknown parameters $\psi$. To estimate these parameters, we employ an EM formulation, drawing inspiration from [31, 29]. While the underlying algorithm follows a similar approach to prior works [57, 40], the formalism has been refined to offer a more precise explanation of the method. The optimization procedure maximizes the expected log-likelihood $\log p_{\psi}(\mathbf{y}|\mathbf{x}_{0})$ under the posterior distribution $p_{\psi}(\mathbf{x}_{0}|\mathbf{y})$
[…]
This objective is unfortunately intractable, because the search quantity $\psi$ appears as a parameter of the distribution used to evaluate the expectation. Therefore, we resort to an EM formulation alternating between an E-step and an M-step. During the E-step, the expectation is evaluated by drawing samples from the approximate posterior $p_{\psi}(\mathbf{x}_{0}|\mathbf{y})$, given fixed RIR parameters $\psi$. The M-step then optimizes the RIR parameters $\psi$ based on the clean speech estimates, maximizing the expected log-likelihood. Both steps are detailed below.
III-C1 E-Step
Given RIR parameters $\psi$, we wish to approximate the posterior speech distribution $p_{\psi}(\mathbf{x}_{0}|\mathbf{y})$ in order to evaluate the expectation in (17). Drawing samples through our posterior sampling procedure explained above in Section III-B would require running the entire diffusion process from $T$ to $T_{\mathrm{min}}$ at each EM iteration, as in [31]. Instead, we follow [29] and evaluate the expectation using the one-step denoised estimate $\hat{\mathbf{x}}_{0}^{\tau}$ through (7) that is obtained at each reverse diffusion step. Formally, the log-likelihood expectation is approximated as
$\mathbb{E}_{p_{\psi}(\mathbf{x}_{0}|\mathbf{y})}\log p_{\psi}(\mathbf{y}|\mathbf{x}_{0})\approx\mathbb{E}_{p(\mathbf{x}_{0}|\mathbf{x}_{\tau})}\log p_{\psi}(\mathbf{y}|\mathbf{x}_{0})\,,$ (18)
and the denoised posterior $p(\mathbf{x}_{0}|\mathbf{x}_{\tau})$ is modelled as a Dirac distribution located at the posterior mean $\hat{\mathbf{x}}_{0}^{\tau}$ [27]
$p(\mathbf{x}_{0}|\mathbf{x}_{\tau})\approx\delta(\hat{\mathbf{x}}_{0}^{\tau})\,.$ (19)
This results in the following one-sample Monte Carlo estimate of the expectation (17)
$\mathbb{E}_{p_{\psi}(\mathbf{x}_{0}|\mathbf{y})}\log p_{\psi}(\mathbf{y}|\mathbf{x}_{0})\approx\log p_{\psi}(\mathbf{y}|\hat{\mathbf{x}}_{0}^{\tau})\,.$ (20)
This effectively integrates the M-step into the reverse diffusion, significantly accelerating inference. Note that in (18), the dependency of $p_{\psi}(\mathbf{x}_{0}|\mathbf{y})$ to $\mathbf{y}$ is no longer explicit. However, this dependency still exists, as we notice that the Tweedie estimate $\hat{\mathbf{x}}_{0}^{\tau}$ is obtained from the current diffusion state $\mathbf{x}_{\tau}$, which has been itself sampled conditionally to $\mathbf{y}$ in (II-B).
III-C2 M-step
Once the expectation in (17) has been evaluated during the E-step, we can proceed to the M-step that will maximize the resulting objective. We add to the expected log-likelihood a regularization term facilitating the RIR parameter search during optimization
$\mathcal{R}(\psi)=\frac{1}{N_{\mathbf{h}}}\rVert S_{\mathrm{comp}}(\mathbf{h}_{\psi})-S_{\text{comp}}(\mathrm{sg}\left[\mathbf{h}_{\psi}\right]+\nu(\tau)\mathbf{v})\ \lVert_{2}^{2},$ (21)
where $\mathbf{h}_{\psi}=\mathcal{A}_{\psi}(\delta)$ is the current time-domain RIR estimate, $\mathbf{v}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is a vector of white Gaussian noise, and $S_{\mathrm{comp}}$ is the magnitude compressed spectrogram, as defined in (10), with $N_{\mathbf{h}}$ time frames. In the right-hand term, the “stopgrad” operator $\mathrm{sg}[\cdot]$ detaches the gradients of $\mathbf{h}_{\psi}$ from the optimization graph and Gaussian noise scaled by $\nu(\tau)$ is added. We show in Appendix -B that this effectively injects multiplicative noise with standard deviation $\nu(\tau)$ in the RIR parameter gradients, taking inspiration from the “regularization by denoising” strategy used, e.g. in [58] for guiding data reconstruction. In our case, this regularization smoothes the RIR parameter optimization landscape by injecting small amounts of stochasticity during the optimization.
The M-step finally consists in maximizing the resulting objective:
[…]
Since all operations, including our reverberation operator $\mathcal{A}_{\psi}(\cdot)$, are differentiable with respect to $\psi$, the M-step can be achieved with gradient-based optimizers, e.g. Adam [53].
During optimization, we further rescale the denoised speech estimate $\hat{\mathbf{x}}_{0}^{\tau}$ so that its root-mean-square power (RMS) matches the average RMS power of clean speech computed on the training set. Using this additional constraint helps lift the indeterminacy when jointly optimizing the speech $\mathbf{x}_{0}$ and RIR parameters $\psi$. This step is included in our ablation study in Section IV-A6. We also found it beneficial to constrain $w_{b}$ and $\alpha_{b}$ within a limited range to stabilize the optimization, specially at early stages. This is achieved by clamping the parameters to predefined minimum and maximum values after every optimization iteration, as specified in Appendix -C1.
IV Experiments and Results